2-(25) Find Pur 102 3:1 341 60/0° 0 Voits Ug 17): 60cosine 3774 volts Z
Answer #1

Ans)

Given vy(t) = 60cos(377t) V

its phasor is

V = 6020 V this is peak phasor as value is peak value as given from above time domain expression

The RMS phasor is

VRMS = - 960 <0°

The load resistance i.e 4 ohm resistor referred to primary is

R = R*(2 = 4* *)2 = 36 9

R = 36 2

Now equivalent circuit referred to primary is

2012 R=36

Current through primary is

60 Ig = VRMS 10 + R 220º 10+ 36 = 0.92220° A

1, = 0.92220 A

As power cannot change in transformer

Pan = P360 = 1,12 * 36 = 0.9222 * 36 = 30.624 W

P4n = 30.624 W